OpenStudy (anonymous):

A ball is kicked upward with an initial velocity of 56 feet per second. The ball's height h (in feet) can be expressed as a function of time t (in seconds) by the equation h = -16t2 + 56t. How much time does the ball take to reach its highest point?

OpenStudy (anonymous):

@ mell6 are in calculus class? to get to the max height, differentiate h, that will be you velocity, v=dh/dt now set this to zero, dh/dt=0, then solve for t, then sub this t to h, that is the max h

OpenStudy (anonymous):

are u in calculus class?

OpenStudy (anonymous):

no but i got it it's like 1.25 i think

OpenStudy (anonymous):

t =1.75 sec sub this to h to get hmm max h= 49 ft? yes or no?

OpenStudy (anonymous):

so you are not in calculus class? h = -16t^2 + 56t. we noticed that this is in the family of parabola y=ax^2 +b or y=-16x^2 +56x plot this on the graph the parabola is opening downward and the vertex is V=-b/2a we get V=-56/2(-16)=1.75 =x=t in sec now sub this to h = -16t2 + 56t we get max h=---?

OpenStudy (anonymous):

@ mell65 let me know if you have question